OpenStudy (anonymous):

WILL FAN & MEDAL! How is the 3888 gotten in this problem? \[\sqrt[4]{3}*\sqrt[5]{2}\] = \[\sqrt[20]{3888}\]

OpenStudy (anonymous):

\[\huge \sqrt[4]{3} = \sqrt[4\times 5]{3^5} = \sqrt[20]{243} \]

OpenStudy (anonymous):

\[\huge \sqrt[5]{3} = \sqrt[5\times 4]{2^4} = \sqrt[20]{16} \]

OpenStudy (anonymous):

\[ \sqrt[20]{16\times 243} = \sqrt[20]{3888} \]

OpenStudy (anonymous):

@Procrasturbator